OpenStudy (anonymous):

simplify : (3x)^2*x^-1 a)3x b)9/x^2 c)9x d)3/x^2

OpenStudy (fibonaccichick666):

so any ideas on how to begin?

OpenStudy (anonymous):

3^2=9 i think it is C 9x

OpenStudy (fibonaccichick666):

3^2=9 is correct, how would you simplify the other parts though?

OpenStudy (fibonaccichick666):

what do you do to the exponents when you multiply like bases?

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

add

OpenStudy (fibonaccichick666):

well yes in this case you would, but that is just adding a negative here

OpenStudy (anonymous):

so it would just b x *9x=9x

OpenStudy (fibonaccichick666):

well you have \(9x^2*x^{-1}\)

OpenStudy (fibonaccichick666):

what you typed that equaled 9x anctually equals \(9x^2\)

OpenStudy (anonymous):

9/x^2

OpenStudy (fibonaccichick666):

how did you get that?

OpenStudy (anonymous):

9x^2*x^-1 2-1=1 9x*x=9x

OpenStudy (fibonaccichick666):

that would be correct, with correct reasoning.

OpenStudy (anonymous):

ok thank u

OpenStudy (fibonaccichick666):

but you do not write 9*x*x, that is 9x^2, there is not inbetween step

OpenStudy (anonymous):

so b is the answer

OpenStudy (anonymous):

?

OpenStudy (fibonaccichick666):

yes, but you are writing 9x^2=9x above, you shouldn't do that. It only works for x=1 then

OpenStudy (anonymous):

or c

OpenStudy (fibonaccichick666):

9x*x=9x^2

OpenStudy (fibonaccichick666):

it's 9x like you said

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

simplify: r^2s^4t/r^2st^3 a)s^3/t^2 b)s^3/t c)s^2/t^3 d)s/t